OpenStudy (anonymous):

Fans & Medals!! I really need help! Complete the equations of the system in slope-intercept form. Use a decimal for the slope if necessary. Line 1 x y 0 3 1 5 Line 2 x y –1 1 –2 –1 Line 1: y =___ x +___ Line 2: y =___ x +___

OpenStudy (sleepyjess):

Do you know how to find slope?

OpenStudy (anonymous):

yea m=y2-y1/x2-x1

OpenStudy (sleepyjess):

Ok, so for the first line, plug in those points

OpenStudy (anonymous):

y2-y1/x2-x1= 5-3/1-0

OpenStudy (sleepyjess):

Ok, and that is 2/1

OpenStudy (sleepyjess):

So 2 is the slope for the first one

OpenStudy (sleepyjess):

Now, when x = 0, what is y?

OpenStudy (anonymous):

1?

OpenStudy (sleepyjess):

Look at the table, x y 0 3

OpenStudy (anonymous):

oh srry was looking at the equation your right =D

OpenStudy (sleepyjess):

Ok, so the equation for the first one is y = 2x + 3

OpenStudy (sleepyjess):

Now, do the same thing for the second line

OpenStudy (anonymous):

I got y=-2/4x+4

OpenStudy (sleepyjess):

How are you getting -2/4?

OpenStudy (anonymous):

y2-y1/x2-x1= -1-1/-2-2=-2/-4 then plug it in to y=mx+b

OpenStudy (dumbcow):

check your points again (-1,1) (-2,-1)

OpenStudy (anonymous):

oh lol ok

OpenStudy (anonymous):

ok so this time i got y=2x+4

OpenStudy (anonymous):

hold on noo

OpenStudy (dumbcow):

not quite, intercept is wrong y = mx + b 1 = 2(-1) + b

OpenStudy (anonymous):

yea srry i was looking at another chart thats why i looked at it again

OpenStudy (anonymous):

ok so after that what do I do??

OpenStudy (dumbcow):

solve for b

OpenStudy (anonymous):

Ok i got 3??

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

Sorry I was doing something but that would make it ifintive?